OpenStudy (anonymous):

solve x^2+6x+7=0 using the quadratic formula

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\qquad \Longrightarrow \qquad x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

do you understand how to do it now?

OpenStudy (unklerhaukus):

sure; you just have to match the equation given with the first one i wrote, hence find a,b,c, plug them into the second equation i wrote, and then simplify